Can you help me figure out this problem please?

Find the distance between the points of intersection of the graphs
y = x^2 + 5 and y = 21
Thank you!
-Tiffany

Answer by fcabanski(1391)

The first graph is a parabola. What is its x value when y = 21? That's where it intersects the line y=21.


21 = x^2 +5


16 = x^2 so x = + or - 4.


The points of intersection are (-4, 21) and (4, 21)


The distance between the two points is 8 units.